# Scalable Training of Graph Foundation Models for Atomistic Materials Modeling: A Case Study with HydraGNN

### v2

@@ Abstract
We present our work on developing and training scalable graph foundation models (GFM) using HydraGNN, a multi-headed graph convolutional neural network architecture. HydraGNN expands the boundaries of graph neural network computations in both training scale and data diversity. It abstracts over message passing algorithms, allowing both reproduction of and comparison across algorithmic innovations that define nearest-neighbor convolution in graph neural networks. This work discusses a series of optimizations that have allowed scaling up the GFM training to tens of thousands of GPUs on datasets that consist of hundreds of millions of graphs. Our GFMs use multi-task learning (MTL) to simultaneously learn graph-level and node-level properties of atomistic structures, such as the total energy and atomic forces. Using over 150 million atomistic structures for training, we illustrate the performance of our approach along with the lessons learned on two state-of-the-art United States Department of Energy (US-DOE) supercomputers, namely the Perlmutter petascale system at the National Energy Research Scientific Computing Center and the Frontier exascale system at Oak Ridge National Laboratory. The HydraGNN architecture enables the GFM to achieve near-linear strong scaling performance using more than 2,000 GPUs on Perlmutter and 16,000 GPUs on Frontier. Hyperparameter optimization (HPO) was performed on over 64,000 GPUs on Frontier to select GFM architectures with high accuracy. Early stopping was applied on each GFM architecture for energy awareness in performing such an extreme-scale task. The training of an ensemble of highest-ranked GFM architectures continued until convergence to establish uncertainty quantification (UQ) capabilities with ensemble learning. Our contribution establishes core capabilities for rapidly developing, training, and deploying further GFMs using large-scale computational resources to enable AI-accelerated materials discovery and design.
@@
ML data parallelism, graph neural networks, message passing, large-scale data processing for ML, atomistic materials modeling
@@ I Introduction
Discovery of new materials with desired properties, accurate predictions of a material’s behavior throughout its entire lifespan, and new chemical processes enabling unparalleled control of chemical transformations and transport are crucial to fundamental scientific progress in energy generation, transportation, electronics, and information technology [1]. Machine learning (ML) has shown great potential in accelerating the screening and pre-selection of materials for further experimental testing. In particular, deep learning (DL) models have shown the ability to effectively capture relevant underlying relationships due to the arrangement of atoms of different constituents within an atomistic structure [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. DL models can be trained on the data generated from experiments and/or first-principles calculations and then used to predict the properties of interest for new inputs. The inference takes only a fraction of the time it would otherwise take to run an experiment or a full first-principles calculation while still producing sufficiently accurate results. This drastic reduction in time to predict material properties using atomistic information results in a promising path towards accelerating material discovery and design [13, 14].
@@
However, generating vast volumes of experimental and/or first-principles data is impractical even with sophisticated experimental facilities and powerful supercomputers. Recently, foundation models (FMs) have demonstrated the capability to navigate around the challenge: once pre-trained over a large volume of available open-source data [15], an FM holds the promise to overcome this limitation by providing a jump-start to refined models by fine-tuning on smaller amounts of data for customized applications (also called downstream tasks). Reducing the number of simulations and/or experiments for generating domain-specific training data also drastically reduces the energy costs of developing domain-specific DL models.
@@
While state-of-the-art language-based FMs with a transformer architecture have reached outstanding results in several domains [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], they fail to capture important topological aspects of the atomistic structures. Therefore, alternative DL architectures that are better suited to retain important topological aspects at the atomistic scale need to be be considered for the development of trustworthy FMs for materials using atomistic scale information.
@@
Since atomistic material structures for a generic type of compound can be mapped onto a graph (where atoms can be treated as nodes and interatomic bonds as edges), graph foundation models (GFMs), which are FMs that operate on data structures as graphs, are the candidate of choice for these applications. Currently, GFMs proposed in the literature are developed by training graph neural networks (GNNs) architectures on a sufficiently large and comprehensive dataset for the domain of interest. While a few efforts have already been undertaken to develop GFMs for atomistic materials modeling applications [32, 33, 34, 35], the existing work is still at an incipient stage. Current efforts do not yet ensure their proposed approach achieves trustworthiness (interpreted as simultaneous achievement of accuracy and high confidence).
The work described in this manuscript is the first-of-its- kind large-scale training of GFMs for atomistic materials modeling. We describe our approach to developing trustworthy supervised GFMs for atomistic materials modeling for the simultaneous prediction of energies and atomic forces. Trustworthiness is obtained (i) by performing HPO at extreme scale to identify a sufficiently broad set of architectures capable of reaching the desired accuracy and (ii) by fully training these HPO candidates to obtain high confidence over the predictions using ensemble UQ. The GFMs have been constructed using HydraGNN [36], a fully scalable GNN architecture developed at ORNL. In addition to full scalability [37], HydraGNN offers several other important capabilities: (i) multi-task learning (MTL) for simultaneously predicting multiple properties and stabilizing the training [38], which we leverage for the simultaneous prediction of energies and atomic forces; (ii) an object-oriented design for message passing neural network (MPNN) layers that allows for automated search of the best performing MPNN by treating the choice of the MPNN as a hyperparameter [39]; (iii) invariant and equivariant features that reduce the computational redundancy and time-to-solution [40], therefore saving energy; and (iv) scalable input/output (I/O) data management for efficient DDP on supercomputing facilities [41]. In addition, for this work we have added important capabilities to HydraGNN for the (a) integration of scalable hyperparameter optimization (HPO) to identify the best-performing configurations of hyperparameters in a computationally efficient manner and (b) scalable ensemble UQ to assess the confidence level of the GFM predictions.
Our focus in this paper is more on the high performance computing (HPC) aspect of the study; we illustrate our approach toward scalable data management, scaling the training process, using HPO at scale, and using ensemble UQ techniques. For training and applying such extremely large GFMs, energy consumption is a paramount concern. During the pre-training of the GFM on large volumes of open-source data, we reduced redundant computations and drastically saved energy by using (1) equivariant features to reduce the computational redundancy and (2) early stopping to select the most promising HydraGNN architectures already at very early initial stages of HPO without fully training architectures that are underperforming. Experiments were conducted on two large US-DOE supercomputers: the Perlmutter petascale machine at National Energy Research Scientific Computing Center (NERSC) and the Frontier exascale system at Oak Ridge National Laboratory (ORNL).
@@
The rest of the paper is organized as follows. We discuss the current state of the art and introduce HydraGNN in Section II. In Section III, we discuss our approach toward developing a scalable framework and list the different optimizations for scalable training. We discuss our use of large-scale HPO to develop a trustworthy GFM. Section IV shows the performance of different components of this work: reading large data, scaling the training process, and performing HPO at large scale. We conclude our study and discuss future work in Section V.
@@ II-A GNN training open-source atomistic data
To date, there have been a few approaches proposed in the literature to develop GFMs for atomistic materials modeling. In [32], the authors proposed a multi-modal approach where multiple encoding DL architectures are trained on different types of data representations and describing different types of quantities. The models are aligned to each other through a penalization term in the training loss function that forces latent vectors from each embedding space to coincide. Even if the approach is proposed to develop FMs to accelerate materials design, the datasets used comprise only organic molecules, which allows to cover only a relatively narrow set of natural elements on the periodic table.
@@
In [34], the authors collected open-source datasets that provide labels for different properties of organic molecules. Using such a diverse collection of datasets, a GNN architecture is used for MTL in order to identify embedding spaces that capture meaningful correlations between the different labeled properties, with the promise that such an embedding space would reduce the data requirement on downstream tasks specific to organic chemistry. Since this approach is trained on open-source datasets that describe only organic molecules, this approach is not transferable to inorganic compounds. Moreover, the authors compare the performance of different MPNN layers to construct the GNN architecture by performing computationally inexpensive hyperparameter tuning on small models with few parameters and transfer the use of such hyperparameters to models of much larger scale. Albeit this approach helps limit the computational burden of HPO on large scale GFMs, the best performing configuration of hyperparameters at small scale is not guaranteed to be the best performing configuration of hyperparameters at a larger scale and on a larger set of data, because the conclusions drawn from the HPO study are model and data dependent.
@@
In [42], the authors developed a GFM trained on the Materials Project Trajectories (MPTrj) dataset [43], using an MPNN layer that is capable to model 4-body interactions. As the authors themselves recognize in their conclusions, albeit their approach sheds light onto a promising path towards building effective GFMs for atomistic materials modeling, the impact of their work is limited by the fact that the GFM has a very small number parameters that was deliberately maintained low due to computational limitations, and this limits the expressivity of the GFM.
@@
While not explicitly presented by their developers as GFMs, there have been other models that cover broader sets of elements of the periodic table compared to the approaches mentioned in the previous paragraph. In [33], the authors built a GNN model using MTL for simultaneously predictions of several material properties by training the GNN model on multiple datasets, including Open Catalyst 2020 (OC2020) [44] and Open Catalyst 2022 (OC2022) [45]. However, the approach considers only a single GNN architecture without performing HPO. Moreover, the set of parameters in the GNN model is relatively small, in the order of few millions of parameters, which limits the attainable accuracy on large volumes of data.
@@
In [35], the authors studied the scaling behavior of 2D molecular GNNs under varied settings of depth, width, number of molecules, number of labels, the diversity in dataset, and the architectural choice. The authors showed that supervised pretraining of large GNNs on molecular datasets provides a rich fingerprint embedding, which is useful for 38 downstream tasks. Even if this work very systematically studied the effect of GNN model size over the predictive performance in the pre-training and fine-tuning stage with many and diverse downstream tasks, the work has two important limitations: it only considers 2D graphs and it addressed only organic compounds.
@@
Several uncertainty quantification (UQ) methods have been applied to GNNs [46], including Bayesian GNNs [47], prediction interval methods [48], and deep ensemble methods [49]. Bayesian methods are theoretically rigorous but challenging to scale to high-dimensional data. Prediction interval methods are cost-effective but often require tedious tuning of heuristic parameters. We leverage deep ensemble methods as a compromise between cost and performance to quantify uncertainty in our GFMs.
@@ II-B Scalability and GPU optimization for GNN training
The effect of the specific algorithmic characteristics of GNNs on performance benchmarking has been carried out on GPUs by [50], where the authors noted that GNN training differs significantly from conventional convolutional networks (CNNs) in that only 25% of the execution time is spent on dense and sparse matrix multiplications compared to 50% in CNNs. Moreover, the execution time to process graph samples in GNNs was noted to vary greatly according to the size of the graph (number of nodes and number of edges) of the input data. The studies conducted in this work showed that the majority of the time during GNN training was spent in integer operations, sorting, index selection, reductions, and scatter-gather operations needed for nodal and edge feature updates with message passing. Multi-GPU scaling was reported using up to 4 GPUs, showing about 20-50% strong scaling efficiency between 1 and 4 GPUs. Similar remarks apply to refs. [51, 52, 53, 54], which characterize subdivision of large graphs among processors and parallel aggregation during convolution steps.
These are useful conclusions for optimization of GNN training on large graphs (i.e., with millions of nodes), but need to be re-evaluated for our datasets. Training on large graphs can be highly sensitive to the splitting scheme used to partition the graph into subgraphs and to distribute them among processors. For atomistic materials modeling applications addressed in our work, the graph samples are small (with at most a few hundreds of nodes). For the GNN convolutions specifically, convolution on a batch of samples will have a much more local, block diagonal structure. Throughput should be less sensitive to the choice of molecules per batch.
Using a larger number of GPUs, the developers of the PyTorch framework for DDP showed the benefit of overlapping computation with communication, showing near-linear scaling using up to 256 NVIDIA Tesla V100 GPUs [55]. These preliminary scaling results focused on DDP for training of DL model using a moderate volume of data. Compared to this preliminary studies, our work shows near-linear scaling using 10x more GPUs and using much larger volumes of data, which introduces important challenges in I/O that we addressed to reduce computational bottlenecks and minimize communication overheads. Moreover, compared to this work, our results are generated using GPUs of newer generations, namely NVIDIA A100 installed on NERSC-Perlmutter and AMD MI250x installed on OLCF-Frontier, thereby showing that our scaling efficiency is also transferable across technologies manufactured by different vendors.
@@ II-C HydraGNN
The complexity of the physics and the scale at which atomistic structures must be studied in response to US-DOE needs in materials science makes it compelling to develop GNN capabilities that simultaneously satisfy several important algorithmic and computer science requirements. To effectively respond to the scientific needs of the US-DOE, a GNN architecture must provide (1) capabilities to read and process data from multiple sources simultaneously, (2) flexibility to support diverse DOE-relevant scientific applications, (3) capabilities to scale the training on leadership class supercomputing facilities, (4) portability across heterogeneous computing environments, (5) continuous software maintenance by ensuring support and compatibility with upgraded software dependencies, (6) maintained documentation to support new users across a broad set of international institutions.
While several GNN architectures have been made available as open-source tools to the scientific community in the last few years [56, 57, 58, 59], none of these tools completely satisfies the above requirements. Moreover, including missing capabilities on these well-established GNN libraries requires invasive and laborious modifications for software re-design. These challenges arising from existing GNN implementations motivated our effort in developing HydraGNN [36, 39], our ORNL-branded, scalable, multi-tasking graph neural network architecture. In response to the US-DOE scientific needs, HydraGNN provides:
@@
  • distributed data management techniques to efficiently scale the training of GNN models on millions of data samples using thousands of GPUs
@@
The HydraGNN library uses the Pytorch [61, 62] software for automatic differentiation and the Pytorch Geometric [63, 64] software for message passing. The architectural hyperparameters that determine the HydraGNN model size and complexity can be set in a configuration file to tune the model training and inference process easily. Overall, HydraGNN is developed and maintained as a high-quality software product for large scale training and development of machine learning models [36].

### v3

# Scalable Training of Trustworthy and Energy-Efficient Predictive Graph Foundation Models for Atomistic Materials Modeling: A Case Study with HydraGNN

@@ Abstract
We present our work on developing and training scalable, trustworthy, and energy-efficient predictive graph foundation models (GFMs) using HydraGNN, a multi-headed graph convolutional neural network architecture. HydraGNN expands the boundaries of graph neural network (GNN) computations in both training scale and data diversity. It abstracts over message passing algorithms, allowing both reproduction of and comparison across algorithmic innovations that define nearest-neighbor convolution in GNNs. This work discusses a series of optimizations that have allowed scaling up the GFMs training to tens of thousands of GPUs on datasets that consist of hundreds of millions of graphs. Our GFMs use multi-task learning (MTL) to simultaneously learn graph-level and node-level properties of atomistic structures, such as energy and atomic forces. Using over 154 million atomistic structures for training, we illustrate the performance of our approach along with the lessons learned on two state-of-the-art United States Department of Energy (US-DOE) supercomputers, namely the Perlmutter petascale system at the National Energy Research Scientific Computing Center and the Frontier exascale system at Oak Ridge Leadership Computing Facility. The HydraGNN architecture enables the GFM to achieve near-linear strong scaling performance using more than 2,000 GPUs on Perlmutter and 16,000 GPUs on Frontier. Hyperparameter optimization (HPO) was performed on over 64,000 Graphic Compute Dies (GCDs) on Frontier to select GFM architectures with high accuracy. Early stopping was applied on each GFM architecture for energy awareness in performing such an extreme-scale task. Each HPO trial was ranked based on both accuracy and energy consumption. The training of an ensemble of highest-ranked GFM architectures (selected with judicious balance between accuracy and energy consumption) continued until convergence to establish uncertainty quantification (UQ) capabilities with ensemble learning. Our contributions establish core capabilities for rapidly developing, training, and deploying further GFMs using large-scale computational resources to enable AI-accelerated materials discovery and design in a trustworthy and energy-efficient manner.
@@
Machine Learning, Distributed Data Parallelism, Graph Neural Networks, Message Passing, Graph Foundation Models, Large-Scale Data Processing for Machine Learning, Atomistic Materials Modeling
@@ I Introduction
Discovery of new materials with desired properties, accurate predictions of a material’s behavior throughout its entire lifespan is crucial to fundamental scientific progress in energy generation, transportation, electronics, and information technology [1]. Machine learning (ML) has shown great potential in accelerating the screening and pre-selection of materials for further experimental testing. In particular, deep learning (DL) models have shown the ability to effectively capture relevant underlying relationships due to the arrangement of atoms of different constituents within various atomistic structures [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. DL models can be trained on data generated from experiments and/or first-principles calculations and then used to predict the properties of interest to be used as new inputs. The inference takes only a fraction of the time it would take to run an experiment or a full first-principles calculation while still producing sufficiently accurate results. This drastic reduction in time to predict material properties using atomistic information results in a promising path towards accelerating material discovery and design [13, 14].
@@
However, generating vast volumes of experimental and/or first-principles training data is impractical even with sophisticated experimental facilities and powerful supercomputers. Recently, foundation models (FMs) have demonstrated the capability to navigate around the challenge: once pre-trained over a large volume of available open-source data [15], an FM holds the promise to provide a jump-start to refined models by fine-tuning on significantly smaller amounts of data for customized applications (also called downstream tasks). Reducing the number of simulations and/or experiments for generating domain-specific training data also drastically reduces the energy costs of developing domain-specific DL models.
@@
While state-of-the-art language-based FMs with a transformer architecture have reached promising results in several domains [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], they fail to capture important topological aspects of the atomistic structures. Therefore, alternative DL architectures that are better suited to retain important topological aspects at the atomistic scale need to be considered for the development of trustworthy (interpreted as simultaneously accurate and highly confident) FMs for materials using atomistic scale information.
@@
Since atomistic material structures for a generic type of compound can be mapped onto a graph (where atoms can be treated as nodes and interatomic bonds as edges), graph foundation models (GFMs) [32, 33], which are FMs that operate on data structures as graphs, are the candidate of choice for these applications. Currently, GFMs proposed in the literature are developed by training graph neural network (GNN) architectures on a sufficiently large and comprehensive dataset for the domain of interest. While a few efforts have already been undertaken to develop GFMs for atomistic materials modeling applications [34, 35, 36, 37], the existing work is still at an incipient stage. Current efforts do not yet ensure that their proposed approach achieves trustworthiness. Moreover, serious concerns have been recently raised about the unaffordable amount of energy requested by proliferating AI centers due to computationally intensive tasks of training large FMs on large volumes of data. Therefore, the development of energy-efficient methodologies for a sustainable development of FMs is of paramount importance.
The work described in this manuscript developed scalable and trustworthy supervised GFMs for the simultaneous prediction of energies and atomic forces with careful attention to consider energy consumption. The GFMs have been constructed using HydraGNN [38, 39], a fully scalable GNN architecture developed at ORNL. Experiments were conducted on two large US-DOE supercomputers: the Perlmutter petascale machine at National Energy Research Scientific Computing Center (NERSC) and the Frontier exascale system at Oak Ridge Leadership Computing Facility (OLCF).
@@
The remaining content of the paper is organized as follows. We discuss the current state of the art and introduce HydraGNN in Section II. In Section III, we discuss our approach toward developing a scalable framework and list the different optimizations for scalable training. We discuss our use of large-scale HPO to develop a trustworthy GFM. Section IV shows the performance of different components of this work: reading large data, scaling the training process, performing HPO at large scale, and measuring epistemic uncertainty using an ensemble of HPO trials with high predictive accuracy and low energy costs. We conclude our study and discuss future work in Section V.
@@ II-A GNN training open-source atomistic data
To date, there have been a few approaches proposed in the literature to develop GFMs for atomistic materials modeling. In [34], the authors proposed a multi-modal approach where multiple encoding DL architectures are trained on different types of data representations and describing different physical quantities. The models are aligned to each other through a penalization term in the training loss function that forces latent vectors from each embedding space to coincide. However, the datasets used comprise only organic molecules, which allows them to cover only a relatively small set of natural elements on the periodic table.
@@
In [36], the authors collected open-source datasets that provide labels for different properties of organic molecules. Using such a diverse collection of datasets, a GNN architecture is used for MTL in order to identify embedding spaces that capture meaningful correlations between the different labeled properties, with the promise that such an embedding space would reduce the data requirement on downstream tasks specific to organic chemistry. Since the model is trained on open-source datasets that describe only organic molecules, this approach is not transferable to inorganic compounds. Moreover, the authors compare the performance of different message passing neural network (MPNN) layers to construct the GNN architecture by performing computationally inexpensive hyperparameter tuning on small models with few parameters and transfer the use of such hyperparameters to models of much larger scale. While this approach helps limit the computational burden of HPO on large scale GFMs, the best performing configuration of hyperparameters at small scale is not guaranteed to be the best performing configuration of hyperparameters at a larger scale and on a larger set of data, because the conclusions drawn from the HPO study are model and data dependent.
@@
In [40], the authors developed a GFM trained on the Materials Project Trajectories (MPTrj) dataset [41], using an MPNN layer that is capable of modeling 4-body interactions. As the authors themselves recognize in their conclusions, while the approach sheds light onto a promising path towards building effective GFMs for atomistic materials modeling, the impact of their work is limited by the fact that the GFM has a very small number parameters that was deliberately maintained low due to computational limitations. Moreover, this reduces the expressivity of the GFM.
In [42], the authors propose a new kernel function for GNNs for modeling potential energy surfaces, which exploits chemical species information, and illustrate its efficacy in the context of transferable GFMs by pretraining the GFM on the Open Catalyst 2020 (OC2020) dataset [43] and illustrating its performance on a set of fine-tuning tasks. Since the GFM was pretrained only on the OC2020 dataset, the applicability of this GFM is restricted to inorganic compounds.
@@
While not explicitly presented by their developers as GFMs, there have been other models that cover broader sets of elements of the periodic table compared to the approaches mentioned in the previous paragraph. In [35], the authors built a GNN model using MTL for simultaneous predictions of several material properties by training the GNN model on multiple datasets, including OC2020 [43] and Open Catalyst 2022 (OC2022) [44]. However, the approach considers only a single GNN architecture without performing HPO. Moreover, the set of parameters in the GNN model is relatively small, in the order of few millions of parameters, which limits the attainable accuracy on large volumes of data.
@@
In [37], the authors studied the scaling behavior of 2D molecular GNNs under varied settings of depth, width, number of molecules, number of labels, the diversity in dataset, and architectural choice. The authors showed that supervised pretraining of large GNNs on molecular datasets provides a rich fingerprint embedding, which is useful for 38 downstream tasks. Even if this work very systematically studied the effect of GNN model size over the predictive performance in the pretraining and fine-tuning stage with many and diverse downstream tasks, the work has two important limitations: it only considers 2D graphs and it addressed only organic compounds.
@@
Several UQ methods have been applied to GNNs [45], including Bayesian GNNs [46], prediction interval methods [47], and deep ensemble methods [48]. Bayesian methods are theoretically rigorous but challenging to scale to high-dimensional data. Prediction interval methods are cost-effective but often require tedious tuning of heuristic parameters. For UQ, we use deep ensemble methods as a compromise between cost and performance to quantify uncertainty in our GFMs.
@@ II-B Scalability and GPU optimization for GNN training
The effect of the specific algorithmic characteristics of GNNs on performance benchmarking has been carried out on GPUs [49], where the authors noted that GNN training differs significantly from conventional convolutional networks (CNNs) in that only 25% of the execution time is spent on dense and sparse matrix multiplications compared to 50% in CNNs. Moreover, the execution time to process graph samples in GNNs was noted to vary greatly according to the size of the graph (number of nodes and number of edges) of the input data. The studies conducted in this work showed that the majority of the time during GNN training was spent in integer operations, sorting, index selection, reductions, and scatter-gather operations needed for nodal and edge feature updates with message passing. Multi-GPU scaling was reported using up to 4 GPUs, showing about 20-50% strong scaling efficiency between 1 and 4 GPUs. Similar remarks apply to refs. [50, 51, 52, 53], which characterize subdivision of large graphs among processors and parallel aggregation during convolution steps.
These are useful conclusions for optimization of GNN training on large graphs (i.e., with millions of nodes), but need to be re-evaluated for our datasets. Training on large graphs can be highly sensitive to the splitting scheme used to partition the graph into subgraphs and to distribute them among processors. For atomistic materials modeling applications addressed in our work, the graph samples are small (with at most a few hundreds of nodes). For GNN convolutions in particular, convolution on a batch of samples will have a much more local, block diagonal structure. Therefore, throughput should be less sensitive to the choice of atomistic structures per batch.
Using a larger number of GPUs, the developers of the PyTorch framework for distributed data parallelism (DDP) showed the benefit of overlapping computation with communication, showing near-linear scaling using up to 256 NVIDIA Tesla V100 GPUs [54]. These preliminary scaling results focused on DDP for training of DL model using a moderate volume of data.
@@ II-C HydraGNN
The complexity of the physics and the scale at which atomistic structures must be studied in response to US-DOE needs in materials science makes it compelling to develop GNN capabilities that simultaneously satisfy several important algorithmic and computer science requirements. To effectively respond to the scientific needs of the US-DOE, a GNN architecture must provide: (1) capabilities to read and process data from multiple sources simultaneously, (2) flexibility to support diverse DOE-relevant scientific applications, (3) capabilities to scale the training on leadership class supercomputing facilities, (4) portability across heterogeneous computing environments, (5) continuous software maintenance by ensuring support and compatibility with upgraded software dependencies, (6) maintained documentation to support new users across a broad set of international institutions.
While several GNN architectures have been made available as open-source tools to the scientific community in the last few years [55, 56, 57, 58], none of these tools completely satisfies the above requirements. Moreover, including missing capabilities on these well-established GNN libraries requires invasive and laborious modifications for software re-design. These challenges arising from existing GNN implementations motivated our effort in developing HydraGNN [38, 39]. In response to the US-DOE scientific needs, HydraGNN provides:
@@
  • invariant and equivariant features that reduce computational redundancy and time-to-solution [61], therefore contributing to energy saving
  • scalable input/output (I/O) data management techniques to efficiently scale the training of GNN models on millions of data samples using thousands of GPUs on supercomputing facilities [62]
@@
HydraGNN uses the PyTorch [63, 64] software for automatic differentiation and the PyTorch Geometric [65, 66] software for message passing. The architectural hyperparameters that determine the HydraGNN model size and complexity can be set in a configuration file to tune the model training and inference process easily. Overall, HydraGNN is developed and maintained as a high-quality software product for large scale training and development of ML models [38].